TITLE: Probability of Head in coin flip when coin is flipped two times
QUESTION [13 upvotes]: Probability of getting a head in coin flip is $1/2$. 
If the coin is flipped two times what is the probability of getting a head in either of those attempts?
I think both the coin flips are mutually exclusive events, so the probability would be getting head in attempt $1$ or attempt $2$ which is:
$$P(\text{attempt $1$}) + P(\text{attempt $2$}) =  1/2 + 1/2 = 1$$
$100\%$ probability sounds wrong? What am I doing wrong. If I apply the same logic then probability of getting at least $1$ head in $3$ attempt will be $1/2+1/2+1/2 = 3/2 = 1.5$ which I know for sure is wrong. What do I have mixed up?

REPLY [2 votes]: After flipping the coin two times there are four possible outcomes: HH, HT, TH, TT. 
Three of them have heads: HH, HT, TH.
So the probability of getting heads in two coin flipps is $3/4$.